OpenStudy (anonymous):

Find the first, fourth, and tenth term of the arithmetic sequence described by the given rule: A(n)=6+(n-1)1/5

OpenStudy (dayakar):

if you want to find out first term write n=1 in the equation A(1)= 6+(1-1)1/5 = 6+ 0 =6 find A(4) A(10)